OpenStudy (anonymous):

A farmer is tracking two wild honey bees in his field. He maps the first bee's path back to the hive on the line y=9/7x. The second bee's path follows the line y=-3x+12. Their paths cross at the hive. At what coordinates will the farmer find the hive?

OpenStudy (anonymous):

So basically solve the system of equations y=9/7x y=-3x+12 you can use substitution, take the expression that y equals in the first equation and plug that into the second equation for y (9/7)x=-3x+12 now can you solve for x

OpenStudy (anonymous):

But I need to use the combination(elimination) method

OpenStudy (anonymous):

Okay, multiply the second equation by -1 so that the ys will cancel y=(9/7)x -y=3x-12 now add them together

OpenStudy (anonymous):

Alright, thanks! (:

OpenStudy (anonymous):

Whataya get?

OpenStudy (anonymous):

Does x=-15/42?

OpenStudy (anonymous):

No =p What do you get when you add the equations together?

OpenStudy (anonymous):

30/7x=-12

OpenStudy (anonymous):

Eh close 0=(30/7)x-12 keep things on whatever side of the equation they were on

OpenStudy (anonymous):

12=30/7x x=84/30x?

OpenStudy (anonymous):

whats 12/(30/7)?

OpenStudy (anonymous):

Isn't it 84/30 because you have to do the reciprocal/

OpenStudy (anonymous):

yep and that simplifies to 14/5 now that you know x plug it into one of the original equations to get y

OpenStudy (anonymous):

Whats -12+12/5? lol

OpenStudy (anonymous):

Where did that come from plug it in y=(9/7)*(14/5) y=?

OpenStudy (anonymous):

y=126/35 simplifies to 3.6?

OpenStudy (anonymous):

Yep thats right

OpenStudy (anonymous):

so the coordinate is (14/5,18/5)

OpenStudy (anonymous):

Omg thank you soo much! (:

OpenStudy (anonymous):

You're welcome =D